OpenStudy (anonymous):

5x+6/x2+2xALL OVERx−4/x2+2x need help simplifying

OpenStudy (anonymous):



hero (hero):

You can't finish the last step from the previous problem? \[\frac{a}{b} \div \frac{b}{c} = \frac{a}{b} \times \frac{b}{c} = \frac{a}{c}\]

OpenStudy (anonymous):

so i dont understand still?

OpenStudy (anonymous):

from @Hero equation a=5x+6 and b=x^2+2x. your c=x-4 and d=x^2+2x

OpenStudy (anonymous):

b=d are the same in this case

OpenStudy (anonymous):

you get the answer?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Same -> change - flip Treat the large fraction as a division problem. The numerator will be the first fraction, the denominator will be the second fraction. Keep the first fraction the same change the division to multiplication flip the second fraction (reciprocal)

OpenStudy (anonymous):

A slight adjustment to Hero's equation to correct a minor mistake. \[\frac{a}{b} \div \frac{c}{b} = \frac{a}{b} \times \frac{b}{c} = \frac{a}{c}\]

OpenStudy (anonymous):

do i cross multiply or direct multiply

OpenStudy (anonymous):

First you see if you can cancel. When multiplying fractions, you can cancel any numerator with any equal denominator.

OpenStudy (anonymous):

Then you multiply straight across.

OpenStudy (anonymous):

do the x^2 + 2x's cancel?

OpenStudy (anonymous):

yes they do.

OpenStudy (anonymous):

5x + 6 over x - 4?

OpenStudy (anonymous):

awesome thank you!

OpenStudy (anonymous):

You got it! Sorry I couldn't add that last bit to your last question, I kept getting booted off.